How many strings can be formed in the length of 5 strings from the word DISCRETE?


1
Expert's answer
2022-06-07T14:00:13-0400

There are 7 different letters in the word "DISCRETE":D,I,S,C,R,E, T.

If the repetition is allowed


"7^5=16807"

We assume that the letter E at the sixth place and the letter E at the eigth place are not different.

If the repetition is not allowed


"\dfrac{7!}{(7-5)!}=2520"